Home‎ > ‎Crew 135‎ > ‎

209-GreenHabReport

Officer Poulet to CapCom and to Coordinator Orenstein.

We received the pH / EC probe today. Please advise on measurements to take (frequency etc.) and in which conditions.
 
The heater has been running until 11:30 at 1500W and I turned it back at 17:30. It was a pretty hot day here on Mars.

The pump is operating normally but we had a power outage over night and the readings I got this morning are not representative of anything.

The seeds I planted on Thursday in my small home-made greenhouse are sprouting whereas the one in the big pots are not showing signs of life yet.

Greenhab readings for 5PM today are as follows:
                         Current             Min                 Max
Temperature     63 Degrees F  57 Degrees F 90 Degrees F   
Humidity           18%                  12%                 18%
 
Greenhab readings at 9am today were as follows:
 
Meter            0.02 KwH
Duration        00h35 min  
 
Best regards,
 
Officer Poulet